OpenStudy (anonymous):

you have a mixture of 40g sand and 50g salt what is the total mass of the mixture? please show me how?

OpenStudy (anonymous):

A simple addition of the two masses will provide the total mass.

OpenStudy (anonymous):

The total mass will be 5g + 4g = 9g. Thank you!

OpenStudy (anonymous):

The total mass will be 5g + 4g = 9g. Thank you!

OpenStudy (anonymous):

The total mass will be 5g + 4g = 9g. Thank you!

OpenStudy (anonymous):

The total mass will be 5g + 4g = 9g. Thank you!

OpenStudy (anonymous):

the total mass will be 40g+50g=90g. I think